# Zhenyu Li1,∗ and Tianyi Shang2 1Shandong Academy of Sciences

### v1

@@ 3.2.2 Row-Column Normalization
where $\log\text{sumexp}(\mathbf{Z},\text{dim}=d)$ denotes the log-sum-exp operation along dimension $d$ for numerical stability. This averaging strategy balances row and column constraints simultaneously, preventing the solution from being dominated by either dimension and enabling more stable convergence.
@@ 4.2.1 Comparison with SOTA Approaches
Table 1: Performance comparison with existing SOTA methods. In the table, bold text represents the best result, and underlined text represents the second-best result.
Pitts30k Pitts250k-test MSLS-val MSLS-challenge
Method R@1/5/10 R@1/5/10 R@1//5/10 R@1/5/10
NetVLAD (CVPR 2016) [4] 81.9/91.2/93.7 90.5/96.2/97.4 53.1/66.5/71.1 35.1/47.4/51.7
SFRS (ECCV 2020) [9] 89.4/94.7/95.9 90.7/96.4/97.6 69.2/80.3/83.1 41.5/52.0/56.3
Patch-NetVLAD (CVPR 2021) [11] 88.7/94.5/95.9 79.2/80.3/83.1 48.1/57.6/60.5
CosPlace (CVPR 2022) [5] 88.5/94.5/95.2 92.4/97.2/98.1 82.8/89.7/92.0 61.4/72.0/76.6
MixVPR (WACV 2023) [2] 91.5/95.5/96.3 94.6/98.3/99.0 88.2/93.1/94.3 64.0/75.9/80.6
R2Former (CVPR 2023) [24] 91.1/ 95.2/96.3 93.2/97.5/98.3 89.7/95.0/96.2 73.0/85.9/88.8
EigenPlaces (ICCV 2023) [6] 92.5/96.8/97.6 94.1/98.0/98.7 89.1/93.8/95.0 67.4/77.1/81.7
SelaVPR (ICLR 2024) [15] 92.8/96.8/97.7 95.7/98.8/99.2 90.8/96.4/97.2 73.5/87.5/90.6
CricaVPR (CVPR 2024) [14] 94.9/97.3/98.2 95.6/98.9/99.5 90.0/95.4/96.4 69.0/82.1/85.7
TransVPR (CVPR 2022) [19] 89.0/94.9/96.2 86.8/91.2/92.4 63.9/74.0/77.5
SALAD (CVPR 2024) [12] 92.4/96.3/97.4 95.1/98.5/99.1 92.2/96.2/97.0 75.0/88.8/91.3
BoQ (CVPR 2024) [3] 94.5/98.9/— 95.0/98.4/99.1 91.4/94.5/96.1
FoL (AAAI 2025) [18] 94.5/97.4/98.2 97.0/99.2/99.5 93.5/96.9/97.6 80.0/90.9/93.0
Pair-VPR (RAL 2025) [10] 95.4/97.5/98.0 95.4/97.3/97.7 81.7/90.2/91.3
SemVPR (ICCV 2025) [23] 94.2/97.7/98.5 89.7/94.7/95.9
$A^{2}$GC 95.6/99.3/99.8 97.3/99.3/99.7 93.6/97.5/97.9 80.6/90.9/92.5
@@ 4.2.2 Ablation Study
Table 6: Experimental evaluate for different input image sizes with the DINOV2_ViTb backbone.

### v2

@@ 3.2.2 Row-Column Normalization
where $\log\text{sumexp}(\mathbf{Z},\text{dim}=d)$ denotes the log-sum-exp operation along dimension $d$ for numerical stability. This averaging strategy simultaneously balances row and column constraints, preventing the solution from being dominated by either dimension and enabling more stable convergence.
@@ 4.2.1 Comparison with SOTA Approaches
Table 1: Performance comparison with existing SOTA methods. Bold is the best result, underline is the second-best result, and * indicates a two-stage re-ranking method.
Pitts30k Pitts250k MSLS-V MSLS-C
Method Desc. size Recall@1/5/10 Recall@1/5/10 Recall@1//5/10 Recall@1/5/10
NetVLAD CVPR’2016[4] 32768 81.9/91.2/93.7 90.5/96.2/97.4 53.1/66.5/71.1 35.1/47.4/51.7
SFRS ECCV’2020[9] 4096 89.4/94.7/95.9 90.7/96.4/97.6 69.2/80.3/83.1 41.5/52.0/56.3
PatchNetVLAD CVPR’2021[11] 4096 88.7/94.5/95.9 79.2/80.3/83.1 48.1/57.6/60.5
CosPlace CVPR’2022 [5] 2048 88.5/94.5/95.2 92.4/97.2/98.1 82.8/89.7/92.0 61.4/72.0/76.6
MixVPR WACV’2023[2] 4096 91.5/95.5/96.3 94.6/98.3/99.0 88.2/93.1/94.3 64.0/75.9/80.6
R2Former CVPR’2023[24] 256 * 91.1/ 95.2/96.3 93.2/97.5/98.3 89.7/95.0/96.2 73.0/85.9/88.8
EigenPlaces ICCV’2023[6] 2048 92.5/96.8/97.6 94.1/98.0/98.7 89.1/93.8/95.0 67.4/77.1/81.7
SelaVPR ICLR’2024[15] 1024 (ViTg) * 92.8/96.8/97.7 95.7/98.8/99.2 90.8/96.4/97.2 73.5/87.5/90.6
CricaVPR CVPR’2024[14] 4096 (ViTb) * 94.9/97.3/98.2 95.6/98.9/99.5 90.0/95.4/96.4 69.0/82.1/85.7
TransVPR CVPR’2022[19] 256 * 89.0/94.9/96.2 86.8/91.2/92.4 63.9/74.0/77.5
SALAD CVPR’2024[12] 8448 (ViTb) 92.4/96.3/97.4 95.1/98.5/99.1 92.2/96.2/97.0 75.0/88.8/91.3
BoQ CVPR’2024[3] 4096 94.5/98.9/— 95.0/98.4/99.1 91.4/94.5/96.1
FoL AAAI’2025[18] 8192 * 94.5/97.4/98.2 97.0/99.2/99.5 93.5/96.9/97.6 80.0/90.9/93.0
Pair-VPR RAL’2025[10] 512 (ViTg) * 95.4/97.5/98.0 95.4/97.3/97.7 81.7/90.2/91.3
SemVPR ICCV’2025[23] 1024 (ViTl) 94.2/97.7/98.5 89.7/94.7/95.9
[…]
@@ 4.2.2 Ablation Study
Table 6: Experimental evaluation for different input image sizes with the DINOV2_ViTb backbone.